Starting From Rest,a fan takes 5 seconds to attain the maximum speed of  400rpm. Assuming constant acceleration,  find the time taken by the fan in attaining  one eighth of it's  maximum speed.

Grade:9

1 Answers

GAMING WITH KRISH
5 years ago
Let the angular  acceleration be alpha. According to the question, 
400rpm=0+5*(alpha )
Let t be the time taken in attaining  the speed of 50rpm which is one eighth of the maximum speed 
Then,  50rpm=0+(alpha)t
Dividing 1st equation by second equation, 
8=5÷(t)
=> t=1.6 seconds
 
Then, 200rpm=0+(alpha)t